OpenStudy (anonymous):

A square pool has an area of 292 square feet. To the nearest tenth, what is the length of one side of the pool?

OpenStudy (anonymous):

\[\sqrt{292}\]

OpenStudy (anonymous):

17.088?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

and put that to the nearest tenth?

OpenStudy (anonymous):

@zimmah ?

OpenStudy (anonymous):

17.1?

OpenStudy (anonymous):

yes thats ur answer!

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Nearest tenth, is first digit after the decimal point, like 17.1

OpenStudy (anonymous):

thx so much

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

can u help help me with one more ?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

A square tile has an area of 148 in2. To the nearest tenth, what is the length of 1 side of the tile?

OpenStudy (anonymous):

That's pretty much the same question as the square pool one, I think you can do this.

OpenStudy (anonymous):

\[\sqrt{148.}\]

OpenStudy (anonymous):

\[\sqrt{148.2}\]

OpenStudy (anonymous):

youcan do it from there

OpenStudy (anonymous):

k

OpenStudy (anonymous):

No it's just 148 in^2, not 148.2

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

oh ok sry

OpenStudy (anonymous):

\[\sqrt{148}\]

OpenStudy (anonymous):

@Lover9 im lost

OpenStudy (anonymous):

alright so do you have a calc?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok do the square root of 148

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

12.165525060596439

OpenStudy (anonymous):

ok now put that into tenths

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

12.1

OpenStudy (anonymous):

or 12.2

OpenStudy (anonymous):

good

OpenStudy (anonymous):

so how long is one side?

OpenStudy (anonymous):

12.1? um like wat do u mean?

OpenStudy (anonymous):

well the question askes how long is one side...so one lide is the length of 12.2 right?

OpenStudy (anonymous):

side*

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok then theres your answer

OpenStudy (anonymous):

lol thx

OpenStudy (anonymous):

any more questions? and np :)

OpenStudy (anonymous):

so its 12.1 yea?

OpenStudy (anonymous):

no 12.2

OpenStudy (anonymous):

yea i forgot the 2 lol

OpenStudy (anonymous):

A square-shaped yard has an area of 114 yd2. To the nearest tenth, what is the length of 1 side of the yard?

OpenStudy (anonymous):

12.1655 is the real one ok? so move the numbers over and you will have 12.2

OpenStudy (anonymous):

oh k

OpenStudy (anonymous):

try doing this yourself :)

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

ill start you off \[\sqrt{144}\]

OpenStudy (anonymous):

this is simple times tables

OpenStudy (anonymous):

10.677078252031311

OpenStudy (anonymous):

10.7?

OpenStudy (anonymous):

uhhh

OpenStudy (anonymous):

\[12\times12\]

OpenStudy (anonymous):

144

OpenStudy (anonymous):

yup theres your answer

OpenStudy (anonymous):

144 ?

OpenStudy (anonymous):

its 12

OpenStudy (anonymous):

no its 12 srry lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

any more questions?

OpenStudy (anonymous):

one more

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

Alex wants to put a fence around a square garden. The garden has an area of 784 ft2. What is the perimeter of the garden?

OpenStudy (anonymous):

does it say the length of one side?

OpenStudy (anonymous):

alright its 112

OpenStudy (anonymous):

Perimeter=112

OpenStudy (anonymous):

wanna fan me?

OpenStudy (anonymous):

i did